Find an equation of the line passing through the given points. Use function notation to write the equation.

​(−2​,−2​) and ​(−4​,−3​)


1
Expert's answer
2022-02-14T16:34:13-0500

The equation of a straight line passing through two points "A(x_1,y_1) \text{~and~} B(x_2,y_2)" is

"\dfrac{y - y_1}{y_2 - y_1} = \dfrac{x - x_1}{x_2 - x_1}".


Here, "A = (-2,-2), B = (-4,-3)". Hence


"\begin{aligned}
\dfrac{y + 2}{-3+2}&= \dfrac{x + 2}{-4+2}\\
2(y+2)&=x+2\\
2y+4&=x+2\\
2y&= x - 2\\
y& = \dfrac{x-2}{2}
\end{aligned}"

The function notation of the straight line is "f(x) = \dfrac{x-2}{2}."